OpenStudy (anonymous):

One number is 2 more than 5 times another. Their product is 72. Find the numbers.

OpenStudy (anonymous):

can somebody show me how to set the equation?

OpenStudy (anonymous):

xy=72 x=5y+2

OpenStudy (anonymous):

a*b=72? what about the other one. 2+5a=b?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so x=-2/5?

OpenStudy (anonymous):

i mean a =-2/5

OpenStudy (anonymous):

try it find b then check

OpenStudy (anonymous):

i got b = 144/5 somehow lol

OpenStudy (anonymous):

and btw theres 2 sets of values

OpenStudy (anonymous):

you have a quadratic equation so should have two answers right?

OpenStudy (anonymous):

question, what exactly is a quadratic equation? lol

OpenStudy (anonymous):

\[x(5x+2)=72\] \[5x^2+2x-72=0\] etc

OpenStudy (anonymous):

equation where the largest exponent is 2

OpenStudy (anonymous):

something that involves factoring and it HAS TO HAVE an equal sign?

OpenStudy (anonymous):

xy=72 x=5y+2 y(5y+2)=72 5y^2+2y-72=0 y=-4===>x=-18 and y=.... x=....

OpenStudy (anonymous):

it has to have an equal sign to be an equation. it has to have degree two to be quadratic

OpenStudy (anonymous):

at ok, thanks for the explanation!

OpenStudy (anonymous):

factoring is a method of solving. in this case you get \[5x^2+2x-72=(5x-18)(x+4)=0\] \[x=-4,x=\frac{18}{5}\]

OpenStudy (anonymous):

and the other set of numbers would probably be 8 and 9?

OpenStudy (anonymous):

halp?

OpenStudy (anonymous):

x=20 y=18/5

OpenStudy (anonymous):

hey satellite, i just wanted to let you know i input your answers and they told me i was incorrect, so i went with ghass and they were right :P thanks for your help though.